OpenStudy (anonymous):

Perform the indicated operations. y^4-16y+2 / y^2+4/5

OpenStudy (unklerhaukus):

\[\frac{y^4-16y+2 }{y^2+4/5}\] ?

OpenStudy (anonymous):

thanks

OpenStudy (unklerhaukus):

is that your question?

OpenStudy (unklerhaukus):

or is it \[y^4-16y+\frac{2}{ y^2}+\frac45\]? or even \[\frac{\frac{y^4-16y+2}{ y^2+4}}5\]?

OpenStudy (anonymous):

not really but still thanks for the help

OpenStudy (unklerhaukus):

is it one of them?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the answer are either 5(y-4) 5(y+4) 5(y-2)